OpenStudy (anonymous):

A sphere of mass 50gm and radius 10cm rolls without slipping with a velocity of 5cm/s. Its total kinetic energy in ergs is?

OpenStudy (anonymous):

KE = translational KE + rotational KE. So translational KE is 1/2mv^2. Rotational KE is 1/2Iw^2. I is the moment of inertia for a solid sphere.

OpenStudy (anonymous):

I think it may need a bit more...maybe not but oh well there are some ideas missing here...the moment of inertia of a sphere is I=mr^2 and the angular velocity is going to be the 5cm/s divided by the radius. Those are going to be need to calculate better...I hope I wasnt going too deep http://mathphystut.tripod.com

OpenStudy (anonymous):

the moment of inertia of a sphere is I=2/5mr^2 you might wanna check that

OpenStudy (anonymous):

ooo...nice call...you a head cracker deeprony

OpenStudy (anonymous):

thanx for the heads up...i dont know how I messed that one up

OpenStudy (anonymous):

no worries. happens to the best of all...=)

OpenStudy (anonymous):

hey can you help me with this one, If (2,-8) is the coordinates of an end point of a focal chord of y^2=4ax, then the co-ordinates of the other end are?